While standing stationary and writing on a patient’s chart, a doctor drops her pen. What can we say about the pen’s motion after she drops it? Ignore air resistance.

A. Its velocity is constant.

B. The magnitude of its velocity is decreasing.

C. The magnitude of its acceleration is increasing.

D. Its acceleration is constant.

Janelle throws a 0.20 kg apple directly upwards with initial speed 16 m/s. Assuming we can ignore air resistance, what will be the height of the apple above the ground after 3.0 seconds? Assume the magnitude of the acceleration due to gravity is g = 9.8 m/s2.

A. 48 m

B. 19 m

C. 3.9 m

D. 1.4 m

A position vector has an x-component equal to 3.2 m and points at an angle 42 degrees below the negative x-axis. What is the magnitude of this vector?

A. 4.3 m

B. 3.2 m

C. 4.8 m

D. cannot be determined from the information given
Answer #1

1) D. Its acceleration is constant.

because, the pen is in free fall.

2) C. 3.9 m

use,

h = voy*t + (1/2)*(-g)*t^2

= 16*3 + (1/2)*(-9.8)*3^2

h = 3.9 m


3) A. 4.3 m

let A is the magnitude of the actual vector.
Ax = 3.2 m

we know,
Ax = A*cos(42)

==> A = Ax/cos(42)

= 3.2/cos(42)

= 4.3